Hi,
Could you please help me to understand the below question?
Point A corresponds to -3 and point B corresponds to 5. What number corresponds to the point that is three times as far from A as it is from B?
Thanks!

Answer by stanbon(75887)

Point A corresponds to -3 and point B corresponds to 5. What number corresponds to the point that is three times as far from A as it is from B?
----
# of units from A to B: 5--3 = 8
----
The point you want is (3/4) of the distance from A to B:
---
Ans: -3 + (3/4)8 = -3+6 = 3
==================================
Cheers,
Stan H.